Find the indefinite integral. (Remember to use absolute values where appropriate. Use C for the constant of integration.)

x2
x − 6

dx
Answer #1

given x2 da ) 2-6 Solution -- I= 73 dre a-6 I = سا x² 36 +36 da x-6 L = s (2_) (x +6) +36 da x-6 36 I = x+6 + ) dx dx a-6 $(x